OpenStudy (anonymous):

Using the formula P = 2(a + b), find a when P = 15 and b = 5. 2.5 5 8 10

OpenStudy (anonymous):

given p=15 and b=5 we just have to replace the value of p and b in that formula so now we get p=2(a+b) =>15=2a+2.5 =>15-10=2a =>5=2a =>2a=5 =>a=5/2 a=2.5

OpenStudy (texaschic101):

just sub 15 in for p and 5 in for b and what do you get ?